ABET CRITERION (E, K) 7-14 A single-phase fluid reservoir is described by three equal gridblocks as shown in Fig. 7-16. The reservoir is horizontal and has homogeneous and isotropic rock properties, k-270 md and 0.27. Gridblock dimensions are Δ-400 ft, 4y-650 ft, and h = 60 ft. Reservoir fluid properties are B = 1 RBSTB and μ = 1 cp. The reservoir left w boundary is kept constant at 3000 psia, and the reservoir right boundary is kept at a pressure gradient of -0.2 psi/ft. Two 7-in vertical wells were drilled at the center of gridblocks 1 and 3. The well in gridblock I injects 300 STB/D of fluid and the well in gridblock 3 produces 600 STB/D of fluid. Both wells have zero skin. Assume that the reservoir rock and fluid are incompressible. Find the pressure distribution in the reservoir es 2 sc3 -600 STB/D qsc1-300 STB/D 60 ft A,=3000 psia 400 ft 400 ft Fig. 7-16. Discretized ID reservoir in Exercise 7-14.
Answer #1

Δx = 400ft; Δy = 650 ft; h = 60 ft; k = 270 md; φ = 0.27; B = 1 RB/STB; µ = 1 cp; Pbw = 3000 psia;

dp/dx|bE = -0.2 psi/ft

Solution:

The gridlocks have same dimensions and properties. So T1,2 = T2,3 = Tx

Tx = βc Axkx / (µBΔx) where βc is transmissibility conversion factor = 0.001127

Tx = 0.001127 * (650*60)*270/(1*1*400) = 29.6683 STB/D-psi

Tx = 29.6683 STB/D-psi

qsc1 = qsc2 = qsc3 = 0

For reservoir 1,

qsc bw-1 = [βc Axkx / (µBΔx/2)]1 [(pbw - p1 ) - γ( Zbw - Z1 )]

= [0.001127*650*60*270/(1*1*400/2)][(3000-p1) -  γ*0] = 59.3366(3000-p1) STB/D

qsc bw-1 = 59.3366(3000-p1) STB/D

For reservoir 3,

  qsc bE-3 = [βc Axkx / (µB)]3 [dp/dx|bE - γ(dZ/dx|bE )]

= [0.001127*650*60*270/(1*1)] [-0.2-(γ*0)] = -2373.462 STB/D

qsc bE-3 = -2373.462 STB/D

The general flow equation for gridblock 1,

T1,2(p2-p1)+qsc bw-1+qsc1 = 0

  29.6683(p2-p1) + 59.3366(3000-p1) + 0 = 0

29.6683p2 - 29.6683p1 + 178009.8 - 59.3366p1 = 0

-89.0049p1 + 29.6683p2 + 178009.8 = 0 ----------------------------------------------- 1

For gridblock 2,

  T1,2(p1-p2)+T2,3(p3-p2)+qsc2 = 0

  29.6683(p1-p2) + 29.6683(p3-p2) +0 = 0

  29.6683p1 - 59.3366p2 + 29.6683p3 = 0  ------------------------------------------------- 2

For gridblock 3,

T2,3(p2-p3)+qsc bE-3+qsc3 = 0

   29.6683(p2-p3) -2373.462 + 0 = 0

29.6683p2 - 29.6683p3 - 2373.462 = 0 -------------------------------------- 3

Solving equation 1, 2 and 3 gives

p1 = 2960 psia;

p2 = 2880.0002 psia;

p3 = 2800.0003 psia;

qsc bw-1 = 59.3366(3000-p1) STB/D

= 59.3366(3000-2960) = 2373.464

qsc bw-1 = 2373.464 STB/D